A software designer is mapping the streets for a new racing game. All of the streets are depicted as either perpendicular or parallel lines. The equation of the lane passing through A and B is -7x + 3y = -21.5. What is the equation of the central street PQ?


1
Expert's answer
2021-07-26T09:16:46-0400

The equation of the lane passing through A & B is "-7x+3y=-21.5"

This could be written as "\frac {7} {3}x -\frac {21.5} {3}"

Compare it with y=mx+c

"m=\frac {7} {3}"


"c=\frac {-21.5} {3}"

Central street PQ will be perpendicular to the lane passing through A & B. Product of slope of perpendicular lines will be"-1"

"m_1.m_2=-1" so "m_2=\frac {-1} {m_1}"


"m_1=\frac {7} {3}"

Slope of perpendicular line "(m_2)=\frac {-1} {slope of parallel line (m_1)}"

"m_2=\frac {-1} {(\frac {7} {3} )}"


"m_2=\frac {-3} {7}"

Equation of line is "y=mx+c"

"y=\frac {-3} {7} x+c"

"7y+3x=7c"

PQ is passing through (7,6),so;

"7(6)+3(7)=7c"

"c=\frac {(41+21)} {7}"

"c=9"

So, Equation of Central street PQ is 

"3x+7y=7c"

"3x+7y=7(9)"

"3x+7y=63"